Jordan Company produces a single product. The projected income statement for the coming month,

based on sales of 20,000 units is:

Sales

$200,000

Less: Variable costs 140,000

Contribution margin $ 60,000

Less: Fixed costs

45,000

Operating income

$ 15,000

Instructions:

1. Compute the unit contribution margin and the number of units that must be sold to break even.

Suppose that 3,000 units are sold above the break-even point. What is the profit?

2. Compute the contribution margin ratio and the break-even point in dollars. Suppose that revenues

are $20,000 greater than expected. What would the total profit be?

3. Compute the margin of safety in sales revenue.

4. Compute the operating leverage.

5. How many units must be sold to earn a profit equal to 10 percent of sales?

6. Assume the income tax rate is 40 percent. How many units must be sold to earn an after-tax profit

of $18,000?
Answer #1

1. Unit contribution margin = 60000/20000 = 3 per unit

Units to be sold for break even = 45000/3 = 15000 units

2. Contribution margin ratio = 3/10 = .30 = 30%

Break even point in dollars = 15000*10 = 150000

Sales Value 170000
Variable cost (17000*7) 119000
Contribution 51000
Fixed Cost 45000
Profit 6000

3. Margin of safety in sales revenue = (200000-150000)*100/200000 = 25%

4. Operating Leverage = 60000/15000 = 4

5. Let x units be sols to earn a profit equal to 10% of sales

sales value 10x
variable cost 7x
contribution 3x
fixed cost 45000
profit (10% of 10x) x

3x - 45000 = x

2x = 45000, x = 22500

6. Let x units be sols to earn a profit of 18000 after income tax of 40%

sales value 20x
variable cost 17x
contribution 3x
fixed cost 45000
profit before tax 3x - 45000
income tax (3x-45000)*40%
profit after tax (3x-45000)*60%

(3x-45000)*60% =18000

1.8x-27000 = 18000

1.8x = 45000

x = 25000